OpenStudy (anonymous):

which number below is a multiple of 3 ? a 222 b 235 c 47 d 239. please show me how to get this answer

OpenStudy (jdoe0001):

a number that's multiple of "3" you'd be able to ADD THE DIGITS and the sum will be divisible by 3

OpenStudy (anonymous):

222

OpenStudy (anonymous):

\[222\div3=74 \]

OpenStudy (texaschic101):

A multiple of 3 is any number divisible by 3

OpenStudy (jdoe0001):

let's say is 239 divisible by 3? let's see 2+ 3+ 9 = 14 can we divide 14 by 3? well no dice, so no -------------------------------------------- let's say it's 54363 divisible by 3? let's see 5 + 4 + 3 + 6 + 3 = 21 can we divide 21 by 3? 21/3 = 7, yes :) so 54363 is divisible by 3 54363/3 = 18121

OpenStudy (texaschic101):

so divide the numbers by 3 and if the answer comes out to be a whole number, it is a multiple of 3.

OpenStudy (texaschic101):

great job explaining jdoe :)

OpenStudy (jdoe0001):

hehe

OpenStudy (amistre64):

what is a number would you have to add up the terms .... and iterate it 3 times, in order to determine if its divisible by 3? ... say the last bit get you down to the 1-12 range of the multiplication table

OpenStudy (amistre64):

wondering if that even makes sense .... its well developed in my head :)

OpenStudy (amistre64):

 such that c is obviously divisible by 3